please help me solve this equation,
The perimeter of a rectangle is 68cm. the perimeter of a right triangle is 60cm. Find the length of the hypotenuse of the triangle. Hypotenuse is 3y-4.
2x+2y=68
???
Charlie
texbook ISBN 0-669-43394-2 Sec. 3.3(Problem solving using linear equations)

Answer by venugopalramana(3286)

please help me solve this equation,
The perimeter of a rectangle is 68cm....WHAT IS THIS PROBLEM?
-----------------------------------------------------
IS THE ONE BELOW ANOTHER PROBLEM??ASSUMING SO
the perimeter of a right triangle is 60cm. Find the length of the hypotenuse of the triangle. Hypotenuse is 3y-4.
LET ONE SIDE = X
PERIMETER = 60
HENCE THIRD SIDE = [60-X)-(3Y-4)]
USING PYTHOGARUS THEOREM
(3Y-4)^2 = X^2+[(60-X)-(3Y-4)]^2=X^2+(60-X)^2+(3Y-4)^2-2(60-X)(3Y-4)
THE GIVEN PROBLEM LEADS TO 2 UNKNOWNS AND ONLY ONE EQN.SO IT IS NOT SOLVABLE.
PLEASE CHECK THE PROBLEM AND COME BACK